# Estimating Fold Changes from Partially Observed Outcomes with Applications in Microbial Metagenomics

### v1

@@ 3.1 Estimation under $H_{A}$
We estimate $\beta$ (our target of inference) and $\mathbf{z}$ (a nuisance parameter) using an estimating equations approach motivated by a Poisson likelihood with mean function (14). To address issues of infinite or undefined MLEs resulting from data separation, which commonly occurs in high-dimensional microbiome sequencing data (Bull et al., 2002), we introduce a Firth penalty on $\beta$ derived from a formal equivalence between our model and the multinomial logistic model. We develop a fast algorithm to find the maximizer of the penalized likelihood by efficiently solving a sequence of score equations that converge to the score of the penalized likelihood within a neighbourhood of the global maximizer. As we will see in Sections 4 and 5, our estimator and accompanying hypothesis testing show excellent performance even when data has excess-Poisson variance, and even in small samples. All algorithms are provided as pseudocode in Supporting Information, and an R package implementing the method is available at [GitHub Link].
To motivate our estimator, we first introduce a Poisson log likelihood for our model. Letting $\beta^{j}$ denote the $j$-th column of $\beta$ and $X_{i}$ denote the $i$-th row of $\mathbf{X}$, the Poisson log likelihood is
$l^{\text{Poisson}}(\beta,z)$ $=\sum_{j=1}^{J}\Big{[}\sum_{i=1}^{n}\Big{(}Y_{ij}(X_{i}\beta^{j}+z_{i})-\exp(X_{i}\beta^{j}+z_{i})\Big{)}\Big{]}.$ (6)
@@
$l^{\text{profile}}(\beta)$ $=\sum_{j=1}^{J}\Big{[}\sum_{i=1}^{n}\Big{(}Y_{ij}\log\frac{\exp(X_{i}\beta^{j})}{\sum_{j^{\prime}=1}^{J}\exp(X_{i}\beta^{j^{\prime}})}-Y_{i\cdot}\frac{\exp(X_{i}\beta^{j})}{\sum_{j^{\prime}=1}^{J}\exp(X_{i}\beta^{j^{\prime}})}\Big{)}\Big{]}+C$ (8)
where $C$ is constant (with respect to $\beta$) and $Y_{i\cdot}:=\sum_{j=1}^{J}Y_{ij}$. Note that the profile likelihood (8) is equal to a multinomial log likelihood with a logistic link (up to a constant). This accords with both known results about the marginal Poisson distribution of multinomial random variables (Birch, 1963), and our observations on identifiability of $\beta$ (only $p\times(J-1)$ parameters can be identified in a multinomial logistic regression of a $J$-dimensional outcome on $p$ regressors).
Unfortunately, the maximizer of $l^{\text{profile}}$ will not be finite if there exists a perfect separating hyperplane (Albert and Anderson, 1984). To both guarantee the existence of an estimate and reduce its bias, we derive the Firth penalty associated with likelihood (8), observing the impact of the nuisance parameters $\mathbf{z}$ on the correct bias-reducing penalty (Kosmidis and Firth, 2011) . Fix $j^{\dagger}\in\{1,\ldots,J\}$ and let $\tilde{\beta}:=[{\beta^{1}}^{T},\ldots,{\beta^{j^{\dagger}-1}}^{T},{\beta^{j^{\dagger}+1}}^{T},\ldots,{\beta^{J}}^{T}]^{T}$, that is, a column vector containing stacked columns of $\beta$ but with $\beta^{j^{\dagger}}$ omitted. Let $\tilde{\mathbf{X}}_{i}=I_{J}^{[-j^{\dagger}]}\otimes X_{i}$, where $I_{J}^{[-j^{\dagger}]}$ indicates a $J\times J$ identity matrix whose $j^{\dagger}$-th column has been removed and $\otimes$ is the Kronecker product, so $\tilde{\mathbf{X}}_{i}\in\mathbbm{R}^{J\times p(J-1)}$ and $\tilde{\mathbf{X}}:=[\tilde{\mathbf{X}}_{1}^{T},\dots,\tilde{\mathbf{X}}_{n}^{T}]^{T}\in\mathbb{R}^{nJ\times p(J-1)}$. Define the fitted values $p_{i}=\exp(\tilde{\mathbf{X}}_{i}\tilde{\beta})\big{/}\mathbf{1}_{J}^{T}\exp(\tilde{\mathbf{X}}_{i}\tilde{\beta})\in\mathbb{R}^{J}$. We can then show that the information matrix associated with (8) is $\tilde{\mathbf{X}}^{T}\mathbf{V}\tilde{\mathbf{X}}$ where $\mathbf{V}$ is a block diagonal matrix with $i$-th $J\times J$ block equal to $Y_{i\cdot}(\text{diag}(\mathbf{p}_{i})-\mathbf{p}_{i}\mathbf{p}_{i}^{T})$ with $\text{diag}(\mathbf{p}_{i})$ indicating a diagonal matrix with diagonal elements equal to $p_{i1},\dots,p_{iJ}$. Finally, we can express the Firth penalized multinomial log likelihood as follows:
$l^{\text{penalized}}=l^{\text{profile}}+\frac{1}{2}\log\big{|}\tilde{\mathbf{X}}^{T}\mathbf{V}\tilde{\mathbf{X}}\big{|},$ (9)
@@
Remarkably, the penalized likelihood (9) is invariant to the choice of $j^{\dagger}$. This is because the Firth penalty is invariant under smooth reparametrizations for full exponential family models such as the multinomial distribution (Jeffreys, 1946; Firth, 1993), and the above parameterization implies the constraint $g(\beta_{k})=\beta_{kj^{\dagger}}=0$. Thus, we can estimate $\beta$ under the “convenience constraint” $\beta_{kj^{\dagger}}=0$ and impose the desired identifiability constraint post-hoc by substracting the relevant $g(\tilde{\beta}_{k})$ from each element of $\tilde{\beta}$. In practice we find that choosing $j^{\dagger}$ to maximize $\sum_{i}\mathbf{1}_{[Y_{ij^{\dagger}}>0]}$ improves the speed of optimization.
To find $\arg\max_{\tilde{\beta}}l^{\text{penalized}}$, we apply the iterative maximum likelihood algorithm of Kosmidis and Firth (2011), which maximizes (9) by iteratively solving multinomial score equations in augmented data $\mathbf{Y}^{(t)}$. Specifically, $Y^{(t)}_{ij}=Y_{ij}+h_{ij}/2$ where $h_{ij}$ is the $(i-1)\times J+j$-th diagonal element of the Poisson “hat” matrix $\mathbf{V}_{p}^{1/2}\tilde{\mathbf{G}}(\tilde{\mathbf{G}}^{T}\mathbf{V}_{p}\tilde{\mathbf{G}})^{-1}\tilde{\mathbf{G}}^{T}\mathbf{V}_{p}^{1/2}$ evaluated at current values of $\beta^{(t)}$ for $\mathbf{V}_{p}$ an $nJ\times nJ$ diagonal matrix with $(i-1)\times J+j$-th diagonal element equal to $\exp(X_{i}\beta^{j}+z_{i})$ and $\tilde{\mathbf{G}}=\Big{[}\tilde{\mathbf{X}}:\mathbf{I}_{n}\otimes\mathbf{1}_{J}\Big{]}\in\mathbb{R}^{nJ\times(p(J-1)+n)}$ (i.e., the expanded design matrix for $\beta$, $\tilde{\mathbf{X}}$, side by side with a design matrix for $\mathbf{z}$, $\mathbf{I}_{n}\otimes\mathbf{1}_{J}$). We compute data augmentations under the restriction on $z$ that $\sum_{j}\exp(X_{i}\beta^{j}+z_{i})=Y_{i\cdot}$ by applying (7). Each iteration $t$ of this algorithm is separable in $z,\beta^{1},\ldots,\beta^{J}$, and we therefore solve each iteration using coordinate descent, applying (7) to update $z$ and updating each $\beta^{j}$ with Fisher scoring followed by a line search. When a desired convergence threshold is reached, we impose the desired identifiability constraint by centering by $g(\tilde{\beta}_{k}^{(t_{\text{max}})})$ as described previously.
@@ 3.2 Estimation under $H_{0}$
In Section 4, we present a robust score test for $\beta_{k^{\star}j^{\star}}=0$ under constraint $g(\beta_{k^{\star}})=0$, which requires fitting a model under the restriction $\beta_{k^{\star}j^{\star}}=g(\beta_{k^{\star}})$. We therefore develop an augmented Lagrangian algorithm (Nocedal and Wright, 2006, p. 514) to solve this constrained optimization problem. The augmented Lagrangian function is
$\mathcal{L}(\beta,z;\rho,u)$ $:=-l^{\text{Poisson}}(\beta,z;\mathbf{Y}^{(t_{\text{max}})})+u\left[g\left(\beta_{k^{\star}}\right)-\beta_{k^{\star}j^{\star}}\right]+\frac{\rho}{2}\left[g\left(\beta_{k^{\star}}\right)-\beta_{k^{\star}j^{\star}}\right]^{2}$ (10)
@@
The presence of $g(\beta_{k^{\star}})$ in (10) complicates optimization, as this term depends on an entire row of $\beta$. As a result, updating $\beta$ column by column (as when fitting models under $H_{A}$) often results in slow convergence under $H_{0}$. Hence, we instead minimize (10) via coordinate descent alternating between block updates to $\beta$ via an approximate Newton step and updates to $z$ via equation (7).
@@
where $\vec{e}_{[k^{\star}j^{\star}]}$ is a vector equal to 1 in the position in $\tilde{\beta}$ corresponding to $\beta_{k^{\star}j^{\star}}$ and zero elsewhere. $\nabla^{2}l^{\text{Poisson}}$ is block diagonal and hence inexpensive to invert, and $\rho(\nabla g-\vec{e}_{[k^{\star}j^{\star}]})(\nabla g-\vec{e}_{[k^{\star}j^{\star}]})^{T}$ is a matrix of rank one. We therefore approximate the Hessian of (10) as $-\nabla^{2}l+\rho(\nabla g-\vec{e}_{[k^{\star}j^{\star}]})(\nabla^{T}g-\vec{e}_{[k^{\star}j^{\star}]})$, and compute its inverse as a rank-one update to $-(\nabla^{2}l)^{-1}$ via the Sherman-Morrison identity (Sherman and Morrison, 1950). For additional details, see Supporting Information Section 3.3.

### v2

@@ 3.1 Estimation under $H_{A}$
We estimate $\beta$ (our target of inference; subject to identifiability constraints) and $\mathbf{z}$ (a nuisance parameter) using an estimating equations approach motivated by a Poisson likelihood with mean function (3). To address issues of infinite or undefined MLEs resulting from data separation, which commonly occurs in high-dimensional microbiome sequencing data, we introduce a Firth penalty on $\beta$ derived from a formal equivalence between our model and the multinomial logistic model. We develop a fast algorithm to find the maximizer of the penalized likelihood by efficiently solving a sequence of score equations that converge to the score of the penalized likelihood within a neighbourhood of the global maximizer. Theorem 7 illustrates some advantages of this proposal: the resultant estimator is normally distributed and consistent for the identifiable parameter $\beta_{k}-g(\beta_{k})$ for any distribution of $Y_{ij}$ satisfying (3), with no requirement for $Y_{ij}$ to be Poisson distributed. In addition, Sections 5 and 6 demonstrate that the estimator and accompanying hypothesis testing show excellent performance even when data has excess-Poisson variance, and even in small samples. All algorithms are provided as pseudocode in Supporting Information (Section 11).
To motivate our estimator, we first introduce a Poisson log likelihood for our model without yet considering the identifiability constraints $g(\beta_{k})=0$. Letting $\beta^{j}\in\mathbb{R}^{p\times 1}$ denote the $j$-th column of $\beta$ and $X_{i}\in\mathbb{R}^{p\times 1}$ denote the $i$-th row of $\mathbf{X}$, the Poisson log likelihood is
$l_{n}^{\text{Poisson}}(\beta,z)$ $=\sum_{j=1}^{J}\left[\sum_{i=1}^{n}\Big{(}Y_{ij}(X_{i}^{T}\beta^{j}+z_{i})-\exp(X_{i}^{T}\beta^{j}+z_{i})\Big{)}\right].$ (6)
@@
$l_{n}^{\text{profile}}(\beta)$ $=\sum_{j=1}^{J}\left[\sum_{i=1}^{n}\Big{(}Y_{ij}\log\frac{\exp(X_{i}^{T}\beta^{j})}{\sum_{j^{\prime}=1}^{J}\exp(X_{i}^{T}\beta^{j^{\prime}})}-\big{|}\big{|}Y_{i}\big{|}\big{|}_{1}\frac{\exp(X_{i}^{T}\beta^{j})}{\sum_{j^{\prime}=1}^{J}\exp(X_{i}^{T}\beta^{j^{\prime}})}\Big{)}\right]+C$ (8)
where $C$ is constant (with respect to $\beta$) and $\big{|}\big{|}Y_{i}\big{|}\big{|}_{1}:=\sum_{j=1}^{J}Y_{ij}$. Note that the profile likelihood (8) is equal (up to a constant) to a multinomial log likelihood with a logistic link. This accords with both known results about the marginal Poisson distribution of multinomial random variables (Birch, 1963), and our observations on identifiability of $\beta$ (only $p\times(J-1)$ parameters can be identified in a multinomial logistic regression of a $J$-dimensional outcome on $p$ regressors).
Unfortunately, the maximizer of $l_{n}^{\text{profile}}$ will not be finite if there exists a perfect separating hyperplane (Albert and Anderson, 1984). To both guarantee the existence of an estimate and reduce its bias, we derive the Firth penalty associated with likelihood (8), observing the impact of the nuisance parameters $\mathbf{z}$ on the correct bias-reducing penalty (Kosmidis and Firth, 2011). Fix $j^{\dagger}\in\{1,\ldots,J\}$ and let $\tilde{\beta}:=[{\beta^{1}}^{T},\ldots,{\beta^{j^{\dagger}-1}}^{T},{\beta^{j^{\dagger}+1}}^{T},\ldots,{\beta^{J}}^{T}]^{T}\in\mathbb{R}^{p(J-1)\times 1}$, that is, a column vector containing stacked columns of $\beta$ but with $\beta^{j^{\dagger}}$ omitted. Let $\tilde{\mathbf{X}}_{i}=I_{J}^{[-j^{\dagger}]}\otimes X_{i}^{T}$, where $I_{J}^{[-j^{\dagger}]}$ indicates a $J\times J$ identity matrix whose $j^{\dagger}$-th column has been removed and $\otimes$ is the Kronecker product, so $\tilde{\mathbf{X}}_{i}\in\mathbbm{R}^{J\times p(J-1)}$ and $\tilde{\mathbf{X}}:=[\tilde{\mathbf{X}}_{1}^{T},\dots,\tilde{\mathbf{X}}_{n}^{T}]^{T}\in\mathbb{R}^{nJ\times p(J-1)}$. Define the fitted values $p_{i}=\exp(\tilde{\mathbf{X}}_{i}\tilde{\beta})\big{/}\mathbf{1}_{J}^{T}\exp(\tilde{\mathbf{X}}_{i}\tilde{\beta})\in\mathbb{R}^{J}$. Then, the information matrix associated with (8) is $\tilde{\mathbf{X}}^{T}\mathbf{V}\tilde{\mathbf{X}}$ where $\mathbf{V}$ is a block diagonal matrix with $i$-th $J\times J$ block equal to $\big{|}\big{|}Y_{i}\big{|}\big{|}_{1}(\text{diag}(\mathbf{p}_{i})-\mathbf{p}_{i}\mathbf{p}_{i}^{T})$ with $\text{diag}(\mathbf{p}_{i})$ indicating a diagonal matrix with diagonal elements equal to $p_{i1},\dots,p_{iJ}$ (Bull et al., 2002). Finally, we can express the Firth penalized multinomial log likelihood as follows:
$l_{n}^{\text{penalized}}=l_{n}^{\text{profile}}+\frac{1}{2}\log\big{|}\tilde{\mathbf{X}}^{T}\mathbf{V}\tilde{\mathbf{X}}\big{|},$ (9)
@@
Remarkably, the penalized likelihood (9) is invariant to the choice of $j^{\dagger}$. This is because the Firth penalty is invariant under smooth reparametrizations for full exponential family models such as the multinomial distribution (Jeffreys, 1946; Firth, 1993), and the above parameterization implies the constraint $g(\beta_{k})=\beta_{kj^{\dagger}}=0$. Thus, we can estimate $\beta$ under the “convenience constraint” $\tilde{\beta}_{kj^{\dagger}}=0$ and impose the desired identifiability constraint $g(\beta_{k})=0$ post-hoc by substracting $g(\tilde{\beta}_{k})$ from $\tilde{\beta}_{k}$. In practice we find that choosing $j^{\dagger}$ to maximize $\sum_{i}\mathbf{1}_{[Y_{ij^{\dagger}}>0]}$ improves the speed of optimization.
To find $\arg\max_{\tilde{\beta}}l_{n}^{\text{penalized}}$, we apply the iterative maximum likelihood algorithm of Kosmidis and Firth (2011), which maximizes (9) by iteratively solving multinomial score equations in augmented data $\mathbf{Y}^{(t)}$. Specifically, $Y^{(t)}_{ij}=Y_{ij}+h_{ij}/2$ where $h_{ij}$ is the $(i-1)\times J+j$-th diagonal element of the Poisson “hat” matrix $\mathbf{V}_{p}^{1/2}\tilde{\mathbf{G}}(\tilde{\mathbf{G}}^{T}\mathbf{V}_{p}\tilde{\mathbf{G}})^{-1}\tilde{\mathbf{G}}^{T}\mathbf{V}_{p}^{1/2}$ evaluated at current values of $\beta^{(t)}$ for $\mathbf{V}_{p}$ an $nJ\times nJ$ diagonal matrix with $(i-1)\times J+j$-th diagonal element equal to $\exp(X_{i}^{T}\beta^{j}+z_{i})$ and $\tilde{\mathbf{G}}=\left[\tilde{\mathbf{X}}:\mathbf{I}_{n}\otimes\mathbf{1}_{J}\right]\in\mathbb{R}^{nJ\times(p(J-1)+n)}$ (i.e., the expanded design matrix for $\beta$, $\tilde{\mathbf{X}}$, side by side with a design matrix for $\mathbf{z}$, $\mathbf{I}_{n}\otimes\mathbf{1}_{J}$). We compute data augmentations under the restriction on $z$ that $\sum_{j}\exp(X_{i}^{T}\beta^{j}+z_{i})=\big{|}\big{|}Y_{i}\big{|}\big{|}_{1}$ by applying (7). Each iteration $t$ of this algorithm is separable in $z,\beta^{1},\ldots,\beta^{J}$, and we therefore solve each iteration using coordinate descent, applying (7) to update $z$ and updating each $\beta^{j}$ with Fisher scoring followed by a line search. When a desired convergence threshold is reached, we impose the desired identifiability constraint by centering by $g(\tilde{\beta}_{k}^{(t_{\text{max}})})$ as described previously to obtain maximizing estimates satisfying $g({\beta}_{k})=0$.
@@ 3.2 Estimation under $H_{0}$
In Section 5, we present a robust score test for $\beta_{k^{\star}j^{\star}}=0$ under constraint $g(\beta_{k^{\star}})=0$, which requires fitting a model under the restriction $\beta_{k^{\star}j^{\star}}=g(\beta_{k^{\star}})$. We therefore develop an augmented Lagrangian algorithm (Jorge and Stephen, 2006, p. 514) to solve this constrained optimization problem. The augmented Lagrangian function is
$\mathcal{L}(\beta,z;\rho,u)$ $:=-l_{n}^{\text{Poisson}}(\beta,z;\mathbf{Y}^{(t_{\text{max}})})+u\left[g\left(\beta_{k^{\star}}\right)-\beta_{k^{\star}j^{\star}}\right]+\frac{\rho}{2}\left[g\left(\beta_{k^{\star}}\right)-\beta_{k^{\star}j^{\star}}\right]^{2}$ (10)
@@
The presence of $g(\beta_{k^{\star}})$ in (10) complicates optimization, as this term may depend on an entire row of $\beta$ (e.g., if $g=g_{p}$). As a result, updating $\beta$ column-by-column (as when fitting models under $H_{A}$) often results in slow convergence under $H_{0}$. Hence, we instead minimize (10) via coordinate descent alternating between block updates to $\beta$ via an approximate Newton step and updates to $z$ via equation (7).
@@
where $\vec{e}_{[k^{\star}j^{\star}]}$ is a vector equal to 1 in the position in $\tilde{\beta}$ corresponding to $\beta_{k^{\star}j^{\star}}$ and zero elsewhere. $\nabla^{2}l^{\text{Poisson}}$ is block diagonal and hence inexpensive to invert, and $\rho(\nabla g-\vec{e}_{[k^{\star}j^{\star}]})(\nabla g-\vec{e}_{[k^{\star}j^{\star}]})^{T}$ is a matrix of rank one. We therefore approximate the Hessian of (10) as $-\nabla^{2}l+\rho(\nabla g-\vec{e}_{[k^{\star}j^{\star}]})(\nabla^{T}g-\vec{e}_{[k^{\star}j^{\star}]})$, and compute its inverse as a rank-one update to $-(\nabla^{2}l)^{-1}$ via the Sherman-Morrison identity (Sherman and Morrison, 1950). For additional details, see Supporting Information (Section 11).
@@
4 Asymptotic properties
We now present consistency, normality and efficiency results for our proposed method. These results demonstrate that if mean model (2) holds, our estimator with identifiability constraint $g(\hat{\beta}_{k})=0$ converges to the identifiable parameter $\beta-g(\beta_{k})$ with a normal limiting law. We also give a stronger condition under which our estimator achieves the asymptotic lower bound on variance. We require no parametric distributional assumptions on $\mathbf{Y}$ (e.g., Poisson, zero-inflated Negative Binomial, etc.) for any of the below results. The absence of parametric assumptions is critical because we believe that parametric models cannot fully capture the characteristics of high-throughput sequencing data.
We begin by considering the identifiability constraint $g(\beta_{k})=\beta_{kJ}$, and then generalize these results to more general smooth constraint functions $g$. Under the parametrization $g(\beta_{k})=\beta_{kJ}$, the model we obtain from profile likelihood (8) is a multinomial logistic regression with reference category $J$.
We begin by showing that the maximizer of the population profile likelihood is the same $\beta$ given by the mean model (2), subject to identifiability. Proofs for all results can be found in Supporting Information Section 9.
Assumption 1.
Let $(X,Y,Z)\overset{iid}{\sim}P^{(0)}$ for unknown true data generating mechanism $P^{(0)}$ with bounded support in $\mathbb{R}^{p}\times\mathbb{R}^{J}\times\mathbb{R}.$ Then $P^{(0)}$ satisfies $\log\mathbbm{E}[Y|X,Z;\beta^{(0)}]=Z\mathbf{1}_{J}+{\beta}^{(0)T}X$ for ${\beta}^{(0)}\in\mathbb{R}^{p\times J}$ with $J$-th column equal to $\mathbf{0}_{J}$.
Assumption 2.
$P^{(0)}$ satisfies $\Sigma_{X}:=Cov(X)$ is positive definite.
Lemma 2.
Under Assumptions 4 and 5, $\beta^{(0)}$ is the unique maximizer of the population criterion $M(\beta):=\mathbbm{E}_{P^{(0)}}\hskip 1.42271ptl_{i}^{\text{profile}}(\beta)$.
We then show that our estimator, which maximizes the Firth-penalized sample likelihood, converges to $\beta^{(0)}$. While we consider a random design, it is straightforward to generalize this result to a fixed design using by placing an eigenvalue condition on the design matrix (Chen et al., 1999).
Assumption 3.
${\beta}^{(0)}$ lies in a closed, bounded subset $\mathbf{\Omega}$ of $\mathbb{R}^{p\times J}$
Theorem 1.
[…]
The demonstration of asymptotic normality is simplified by considering the vector-valued parametrization $\tilde{\beta}\in\mathbb{R}^{p(J-1)\times 1}$ of the matrix-valued parameter $\beta\in\mathbb{R}^{p\times(J-1)}$. For $X\in\mathbb{R}^{p}$, we move between the vector notation $\tilde{\beta}$ and matrix notation $\beta\in\mathbb{R}^{p\times(J-1)}$ using the equality $\tilde{X}\tilde{\beta}={\beta}^{T}X$ for $\tilde{X}=I_{J}^{[-J]}\otimes X^{T}\in\mathbb{R}^{J\times p(J-1)}$, noting that we drop the $J$-th column of $\beta$ via $g(\beta_{k})=\beta_{kJ}$.
Theorem 2.
Under Assumptions 46, the MPLE $\sqrt{n}\left(\hat{\tilde{\beta}}^{(n)}-\tilde{\beta}^{(0)}\right)\overset{\mathcal{D}}{\rightarrow}\mathcal{N}(\mathbf{0},\Sigma)$ for positive definite $\Sigma$.
Given the above results for $\hat{\beta}\in\mathbb{R}^{p\times J}$ with $\hat{\beta}_{\cdot J}=\mathbf{0}_{p}$, expanding the above results to generic constraint function $g(\hat{\beta}_{k})=\mathbf{0}_{p}$ only requires the smoothness of $g$. Choosing $g$ to satisfy $g(\mathbf{b}+a)=g(\mathbf{b})+a$ is not necessary for Theorem 7. That said, the identifiability result of Lemma 1 as it relates to estimating the parameter associated with the unobserved data $\mathbf{W}$ means this additional requirement on $g$ is a useful practical restriction.
Theorem 3.
[…]
We conclude this section by showing that our proposed estimator attains the smallest possible asymptotic variance when the variance of the observations is proportional to either Poisson or Multinomial variance. While we do not necessarily believe either variance assumption, this result demonstrates when our proposed estimator is not just asymptotically correct, but also optimal within a desirable class of estimators.
Theorem 4.
Under Assumptions 46 and either $\text{Var}(Y|X,Z;\beta^{(0)})=\phi\times$ $\text{diag}\left\{\mathbbm{E}[Y|X,Z;\beta^{(0)}]\right\}$ or $\text{Var}(Y|X,Z;\beta^{(0)})=\phi\tilde{\mathbf{V}}$ for unknown constant $\phi$, $\hat{\beta}_{n}$ is asymptotically efficient among asymptotically linear estimators of $\beta^{(0)}$.